OpenStudy (anonymous):

What is the solution set for |–2x + 16| – 9 = –3

OpenStudy (saifoo.khan):

|–2x + 16| – 9 = –3 |–2x + 16| = 6

OpenStudy (saifoo.khan):

–2x + 16 = 6 or –2x + 16 = -6

OpenStudy (katrinakaif):

Bring the -9 to the other side by adding it |-2x + 16| = -3 + 9 |-2x + 16| = +6 Substract 16 to both sides |-2x| = -10 |x| = 5

OpenStudy (saifoo.khan):

–2x + 16 = 6 -2x = -10 x = 5 or –2x + 16 = -6 -2x = -22 x = -11